Express A^5-A^4+A^2-4I as linear polynomial in A where A=metrix 3 1

-2 2


1
Expert's answer
2021-01-11T16:09:04-0500

Given matrix is "A=\left [ \begin{matrix}
 3 & 1 \\
 -2 & 2
 \end{matrix}
\ \right ]"

Find the characteristic polynomial equation for the given matrix "A".

A characteristic polynomial equation is given by

"\left | A-\lambda I \right |=0", where "\lambda" is any scalar.

Using the given matrix, the characteristic equation is

"\left | 
 \begin{matrix}
 3-\lambda & 1 \\
 -2 & 2-\lambda 
 \end{matrix}
\ \right |=0"

"\Rightarrow (3-\lambda )(2-\lambda )+2=0"

"\Rightarrow 6-3\lambda -2\lambda +\lambda ^2+2=0"

"\Rightarrow \lambda ^2-5\lambda +8=0"

Caley Hamilitons theorem states that every square matrix satisfies its characteristic equation.

That is, we get "\Rightarrow A ^2-5A +8I=0" ... (1)

Now let us take the given expression "A^5-A^4+A^2-4I" and express it as linear polynomial using the equation (1)

Now "A^5-A^4+A^2-4I=A^3(A^2-5A+8I)+4A^4-8A^3+A^2"

"=A^3(0)+4A^4-8A^3+A^2" ( using (1) )

"=4A^4-8A^3+A^2"

"=4A^2(A^2-5A+8I)+12A^3-31A^2"

"=4A^2(0)+12A^3-31A^2"

"=12A^3-31A^2"

"=12A(A^2-5A+8I)+29A^2"

"=12A(0)+29A^2"

"=29A^2"

"=29(A^2-5A+8I)+145A-232I"

"=29(0)+145A-232I"

"=145A-232I"

Therefore, "A^5-A^4+A^2-4I=145A-232I"